A golfer hit a golf ball at 35° from the ground with a velocity of 43m/s what is the height of the golf ball at the apex of its flight
1
Expert's answer
2020-11-02T09:23:20-0500

We can find the height of the golf ball at the apex of its flight from the kinematic equation:


"v_y^2=v_{0y}^2-2gy_{max},"

here, "v_y=0\ \dfrac{m}{s}" is the velocity of the golf ball at the maximum height, "v_{0y}=v_0sin\theta" is the vertical component of the initial velocity of the golf ball, "v_0=43\ \dfrac{m}{s}" is the initial velocity of the golf ball, "\theta=35^{\circ}" is the launch angle, "y_{max}" is the height of the golf ball at the apex of its flight (or the maximum height attained by the golf ball) and "g = 9.8 \ \dfrac{m}{s^2}" is the acceleration due to gravity.

Then, from this formula we can find "y_{max}":


"0=v_0^2sin^2\theta-2gy_{max},""y_{max}=\dfrac{v_0^2sin^2\theta}{2g} = \dfrac{(43\ \dfrac{m}{s})^2 \cdot sin^2 35^{\circ}}{2\cdot 9.8 \ \dfrac{m}{s^2}} = 31 \ m."

Answer:

"y_{max}=31 \ m."